OpenStudy (anonymous):

Use the following four coordinates to determine which one has a distance of 5 units from point A (2, 6). (7, 5) (6, 7) (2, 5) (5, 2)

OpenStudy (anonymous):

@LanguageEnthusiast I chose answer (6,7)

OpenStudy (anonymous):

\[d= \sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] Do you know how to use that formula?

OpenStudy (anonymous):

I believe so. Let me give it a try

OpenStudy (anonymous):

2,6) (6,7) D= 2-6)2 + (6-7)2 D= (4)2 + (1)2 D= 16 + 1 D= 17? o-o

OpenStudy (anonymous):

Yup, now just keep doing that until you find which one gives you: \[d=5\]

OpenStudy (anonymous):

I think it's (7,5).... I made a grid... and this one was 5 units apart.

OpenStudy (anonymous):

(2,6) (7,5) D=(2-7)2 + (6-5)2 D=(-5)2 + (1)2 D=25 + 1 D= 26

OpenStudy (anonymous):

Not right ._.

OpenStudy (anonymous):

ERMAHGERD

OpenStudy (anonymous):

Try: \[(5,2)\]

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

(2,6) (5,2) D= (2-5)2 + (6-2)2 D= (3)2 + (4)2 D= 9 + 16

OpenStudy (anonymous):

? ;-;

OpenStudy (anonymous):

\[9 + 16 = 25\] \[\sqrt{25}=5\]

OpenStudy (anonymous):

:D Thank you. I always forget that last step.

OpenStudy (anonymous):

No problem, it happens to all of us. :-P

OpenStudy (anonymous):

Help me out with a few more? This will help me get my grade up quite a lot. Math really stresses me out... I over think it.